OpenStudy (anonymous):

Find the standard equation of the hyperbola with vertices(0,+/-6) and asymptote y=2x

OpenStudy (anonymous):

This is a vertical hyperbola, and thus the slopes of the asymptotes are m = a/b. Since a = 6, b = 3. The center is at the origin. The equation is y^2/36 - x^2/9 = 1